### v1

# Refining Neural Activation Patterns for Layer-Level Concept Discovery in Neural Network-Based Receivers

Abstract
Concept discovery in neural networks often targets individual neurons or human-interpretable features, overlooking distributed layer-wide patterns. We study the Neural Activation Pattern (NAP) methodology, which clusters full-layer activation distributions to identify such layer-level concepts. Applied to visual object recognition and radio receiver models, we propose improved normalization, distribution estimation, distance metrics, and varied cluster selection. In the radio receiver model, distinct concepts did not emerge; instead, a continuous activation manifold shaped by Signal-to-Noise Ratio (SNR) was observed—highlighting SNR as a key learned factor, consistent with classical receiver behavior and supporting physical plausibility. Our enhancements to NAP improved in-distribution vs. out-of-distribution separation, suggesting better generalization and indirectly validating clustering quality. These results underscore the importance of clustering design and activation manifolds in interpreting and troubleshooting neural network behavior.
1 Introduction
Neural networks are increasingly used in high-stakes applications (e.g., Jumper et al. (2021))—yet their internal representations remain difficult to interpret (Molnar (2022)). This opacity complicates debugging, trust, and accountability, especially in domains where model behavior must align with physical or safety constraints (Huyen (2022); Adebayo et al. (2020)). With regulatory frameworks like the EU AI Act mandating transparency in AI systems (European Commission (2024, 2019)), interpretable machine learning has become both a technical and societal necessity.
6G is a high-stakes application, projected to reach a \$68.7B global market by 2035 with a CAGR of 76.9% (MarketsAndMarkets (2024)), with the first commercial deployments expected by 2030 (Nokia (2025)). It promises extreme data rates, ultra-low latency, and dense connectivity, driving the need for advances in spectrum use, ultra-massive MIMO, and adaptive signal processing (Nokia Bell Labs (2024)). At the core is the physical layer, which sets the fundamental limits for capacity, efficiency, and reliability. Errors here propagate upward, making its performance critical. Traditional signal processing (Goldsmith (2005); Tse and Viswanath (2005)) is reaching its limits—motivating a shift to data-driven models with better adaptability (Hoydis et al. (2021)).
We explore this shift in the context of data-driven physical-layer receivers, where emerging 6G architectures are expected to replace hand-crafted pipelines with end-to-end learned models (Hoydis et al. (2021)). Architectures like DeepRx (Honkala et al. (2021)) process raw physical-layer radio signals to decode transmitted information under dynamic, noisy channel conditions. Consequently, these models introduce a new challenge: interpretability (Tuononen et al. (2025)). Physical-layer inputs are non-semantic, combining transmitted data with unpredictable environmental effects (e.g., noise, fading, interference), making most standard interpretability methods ineffective (Chergui et al. (2025); Fiandrino et al. (2022)).
To address this, we build on the Neural Activation Pattern (NAP) methodology (Bäuerle et al. (2022)), which clusters full-layer activation distributions to identify distributed, layer-level concepts. We extend NAP with improved normalization, distribution estimation, distance metrics, and varied cluster selection—as outlined in Figure 1—resulting in fewer unclustered samples and more stable clustering, and better out-of-distribution detection accuracy. To interpret the discovered clusters, we compare them with known physical parameters such as Signal-to-Noise Ratio (SNR). While no discrete cluster structure emerged, lower-dimensional embeddings revealed a continuous activation manifold aligned with SNR—suggesting neural receivers implicitly learn representations structured by core physical properties, even without explicit supervision.
Figure 1: Layer-level concept discovery with proposed improvements highlighted in PROPOSED. The top part illustrates the DeepRx model processing received signals to estimate transmitted bits. The bottom part outlines the concept discovery pipeline operating at the layer level.
2 Methodology
We refine the original NAP methodology (Bäuerle et al. (2022); detailed in Appendix A) to improve the stability and generalizability of the clustering results. In addition, we propose a principled approach to interpreting and evaluating clustering in the context of neural network-based systems, with a focus on radio receivers. Our refinements are shown in Figure 2 and described below:
Proposed Data Extraction: Follows the same procedure as in the baseline NAP methodology, where we extract activation outputs from a specified layer $l$ of a neural network $n$ for a given input set $X=\{x_{i}\mid i\in[1,\ldots,N]\}$. This yields a set of activations $A^{l}=\{a^{l}_{i}=n^{l}(x_{i})\}$, where each $a^{l}_{i}$ contains the output feature maps from layer $l$. In convolutional layers, these are structured as per-channel activation matrices, used in subsequent processing.
Proposed Normalization: We rescale the activations of each convolutional channel to mean 0 and standard deviation 1 (known as Z-Score Normalization), making them more suitable for subsequent distribution estimation by removing scale differences across channels and centering the data:
$\phantom{,}\hat{a}^{l}_{i,u}=\frac{a^{l}_{i,u}-\mu^{l}_{u}}{\sigma^{l}_{u}},$ (1)
where $\hat{a}^{l}_{i,u}$ represents the normalized activation, $a^{l}_{i,u}$ is the original activation, $\mu^{l}_{u}$ is the mean, and $\sigma^{l}_{u}$ is the standard deviation of the activation output of channel $u$ in layer $l$ across all data samples in $X$.
Proposed Location Disentanglement: We employ a form of non-parametric density estimation technique, known as Kernel Density Estimator (KDE; Rosenblatt (1956); Parzen (1962)), to aggregate the normalized activations per convolutional channel. For KDE, the estimated Probability Density Function (PDF) for channel $u$ in layer $l$ is given by
$\phantom{,}\hat{f}^{l}_{i,u}(a)=\frac{1}{QH}\sum_{q=1}^{Q}K\left(\frac{a-\hat{a}^{l}_{i,u,q}}{H}\right),$ (2)
where $K()$ is the kernel function (e.g., Gaussian), $H$ is a smoothing parameter called the bandwidth, and $Q$ is the number of activation values on channel $u$ in layer $l$. We use Scott’s Rule (Scott (1992)) for selecting the bandwidth, providing a standard bias-variance tradeoff:
$\phantom{.}H=1/\sqrt[5]{Q}.$ (3)
We sample the estimated PDF (Equation 2) over a Z-score range covering $P\%$ of the normalized activation mass to reduce outlier influence:
$\phantom{,}z_{\text{min}}=\Phi^{-1}\left(\frac{1-\tfrac{P}{100}}{2}\right),\quad z_{\text{max}}=\Phi^{-1}\left(1-\frac{1-\tfrac{P}{100}}{2}\right),$ (4)
where $\Phi^{-1}(\cdot)$ is the Inverse Cumulative Distribution Function (ICDF) of the standard normal distribution. The range is then discretized into $R$ uniformly spaced points for channel $u$ in layer $l$ to produce a fixed-length representation, as follows:
$\phantom{.}\mathbf{\hat{f}}^{l}_{i,u}=\left[\hat{f}^{l}_{i,u}\left(z_{\text{min}}+r\cdot\frac{z_{\text{max}}-z_{\text{min}}}{R-1}\right)\right]_{r=0}^{R-1}.$ (5)
Proposed Clustering Analysis: We cluster the sampled and concatenated estimated density values
$\phantom{,}\mathbf{\hat{g}}^{l}_{i}=\left[\hat{f}^{l}_{i,1,1},\hat{f}^{l}_{i,1,2},\dots,\hat{f}^{l}_{i,1,R},\hat{f}^{l}_{i,2,1},\hat{f}^{l}_{i,2,2},\dots,\hat{f}^{l}_{i,U,R},\right]_{u=1}^{U},$ (6)
where $U$ is the number of channels in layer $l$, using HDBSCAN* with the Jensen-Shannon distance metric (Lin (1991); Endres and Schindelin (2003)), which is symmetric, bounded, and well-suited for comparing probability distributions such as those estimated by KDE in the previous step. The Jensen-Shannon distance $D^{l}_{\text{JS}}$ between activations for inputs $x_{i}$ and $x_{j}$ for all channels in layer $l$ is
$D^{l}_{\text{JS}}(x_{i},x_{j})=\frac{1}{\sqrt{2}}\sqrt{\sum^{U\cdot R}_{k=1}{\left(\mathbf{\hat{g}}^{l}_{i,k}\log\frac{2\cdot\mathbf{\hat{g}}^{l}_{i,k}}{\mathbf{\hat{g}}^{l}_{i,k}+\mathbf{\hat{g}}^{l}_{j,k}}+\mathbf{\hat{g}}^{l}_{j,k}\log\frac{2\cdot\mathbf{\hat{g}}^{l}_{j,k}}{\mathbf{\hat{g}}^{l}_{i,k}+\mathbf{\hat{g}}^{l}_{j,k}}\right)}}$ (7)
and we use this as a distance metric in clustering.
Figure 2: Discovering layer-level concepts by applying NAP methodology with our proposed improvements (underlined): extracting, preprocessing, and clustering activations—followed by interpreting and evaluating the clustering results.
Proposed Post-Clustering Analysis: We evaluate clustering quality using two internal metrics and one application-specific metric. As internal metrics, we use the persistence score (Campello et al. (2013)), which reflects the stability or robustness of a cluster (with higher values indicating more well-defined clusters), and the Density-Based Clustering Validation (DBCV) index (Moulavi et al. (2014)), which is designed to assess density-based clustering algorithms like HDBSCAN by balancing cluster compactness and separation. As an application-specific metric, we use Out-of-Distribution (OOD) detection (see, e.g. Zhu et al. (2025)), assuming that OOD samples fall outside the learned cluster structure. This allows clustering to act as a proxy for distributional mismatch. In this setting clustering serves not just as an explanation method but as a model debugging tool—similar to how Adebayo et al. (2020) frame OOD behavior as test-time contamination. To interpret the resulting clusters, we use the channel parameters of the inputs (e.g., SNR) to summarize and explain structure.
3 Experiments and Results on Visual Object Recognition Model Activations
We applied the NAP methodology using the average amount of activations as the baseline, as suggested by Bäuerle et al. (2022) to be the most suitable choice for extracting a larger number of NAPs. We then compared this baseline with our proposed methodology—outlined in Figure 2 and detailed in Section 2—and additionally studied the effect of HDBSCAN parameters.
3.1 Experimental Setup
The deep neural network-based visual object classification model was based on the ResNet-50 architecture He et al. (2016), using a pretrained weights from TorchVision (Marcel and Rodriguez (2010)). Further implementation and dataset details are provided in Appendix C.
We used our own PyTorch-based implementation of both the NAP baseline and our proposed methodology; implementation details are provided in Appendix B.
3.2 Results
We first evaluated the impact of dataset size using the HDBSCAN parameters recommended by Bäuerle et al. (2022)—specifically, the Leaves of Tree (Leaf) as the cluster selection criteria with a minimum cluster size of 5. As shown in Figure 4, the baseline method resulted in 73–95% of samples remaining unclustered. Our proposed methodology did not mitigate this issue. A high proportion of non-clustered activations is problematic, as it hinders the discovery of meaningful layer-level concepts, reducing interpretability. It also undermines confidence in the clustering structure and impairs downstream tasks such as Out-of-Distribution (OOD) detection, which depend on well-formed, dense regions in activation space.
Notably, Bäuerle et al. (2022) do not report any analysis of unclustered or outlier samples, which we found to be a recurring issue in the visual object recognition setting as well. Further results (Appendix E) show that increasing dataset size leads to more clusters being discovered—consistent with findings in Bäuerle et al. (2022)—but the mean persistence score remains relatively stable.
Next, we investigated the effect of HDBSCAN parameters on clustering outcomes. Figure 4 shows that using Excess of Mass (EoM) as the cluster selection criteria consistently reduces the number of unclustered samples across settings. Smaller minimum cluster sizes (e.g., 3) also lead to fewer unclustered samples. With the updated cluster selection criteria (EoM), our methodology generally clusters more samples—except for Conv5 with a minimum cluster size of 5. The number of clusters behaves differently depending on the cluster selection method: with Leaf, the number of clusters decreases as the minimum cluster size increases, likely due to aggressive pruning of small clusters; with EoM, the number remains more stable, as it prioritizes persistence over strict pruning.
Figure 3: Percentage of clustered and non-clustered samples as a function of sample size and aggregation method with minimum cluster size 5 using ImageNet subset with ResNet-50 model.
Figure 4: Number of non-clustered samples and clusters by aggregation and cluster selection method as function of minimum cluster size using ImageNet subset with ResNet-50 model.
As our primary focus is on analyzing activations from the radio receiver model, we conclude our analysis of the visual object recognition results here and proceed to the next set of experiments.
4 Experiments and Results on Radio Receiver Model Activations
Following the approach used in Section 3, we applied both the baseline and proposed methodologies—illustrated in Figure 2 and described in detail in Section 2 and in Appendix A—to a generated dataset. Our goal was to analyze and summarize the layer-level concepts learned by the deep neural network-based receiver model in terms of underlying channel parameters.
4.1 Experimental Setup
The deep neural network-based receiver model (DeepRx, as detailed in Appendix D) was based on Honkala et al. (2021), adopting the Performer Model from Table 1 of Tuononen et al. (2025), which presents a more minimal version. The model was trained following Honkala et al. (2021). Training and testing data were generated using link-level simulations with the NVIDIA Sionna library (Hoydis et al. (2023)), as detailed in Table 1 in Appendix D. We applied the same NAP baseline and proposed methodology as described in Section 3.1, with implementation details provided in Appendix B.
4.2 Clustering Sensitivity to Parameters and Preprocessing
We first evaluated the impact of dataset size using HDBSCAN with the Leaf cluster selection criterion and a minimum cluster size of 5, as recommended by Bäuerle et al. (2022). As shown in Figure 6, the baseline method resulted in 59–97% of samples remaining unclustered, with the proportion increasing alongside dataset size. Our proposed methodology did not resolve this issue. As in the previous experiment (Section 3.2), more data led to more clusters being discovered (Appendix F), but here the mean persistence score tended to decrease.
We then explored the effect of HDBSCAN parameters. As shown in Figure 6, using the EoM criterion and smaller minimum cluster sizes (e.g., 3 or 5) consistently reduced the number of unclustered samples. Our proposed method generally offered more stable performance, avoiding sharp changes seen in the baseline (e.g., at cluster size 13 on B1-PRE, or high unclustered rates on OUT-POST). Cluster count trends mirrored those in the ImageNet experiment, except for OUT-POST, where both criteria showed decreasing cluster counts with increasing minimum cluster size.
Applying Uniform Manifold Approximation and Projection (UMAP) for dimensionality reduction with standard settings before HDBSCAN (Appendix L) generally increased both the number of clusters and unclustered samples, making interpretation more difficult. Although a few configurations (e.g., proposed method with EoM, minimum cluster size 19, and 15 UMAP dimensions on B1-PRE) reduced unclustered samples to near-zero, these effects were inconsistent across layers and methodologies. We suspect that UMAP may distort the data structure in ways that occasionally aid clustering, but its behavior was too unstable to be useful in practice (Appendix I).
Finally, applying HDBSCAN directly to raw activations—i.e., without aggregation or normalization—left all samples unclustered (see Appendices J and L), likely due to the curse of dimensionality (Bellman (1961)). Adding UMAP for dimensionality reduction in this setup sometimes enabled full clustering—for example, using EoM as the cluster selection criterion with a minimum cluster size of 3 on layer B1-PRE and 15 UMAP dimensions—but results remained erratic and inconsistent across layers and parameters. As such, we do not further consider the "no aggregation and no normalization" methodology or UMAP-based dimensionality reduction in this work.
Figure 5: Percentage of clustered and non-clustered samples as function of sample size and aggregation method with minimum cluster size 5 using link-level simulations with DeepRx model.
Figure 6: Number of non-clustered samples and clusters by aggregation and cluster selection method as function of minimum cluster size using link-level simulations with DeepRx model.
4.3 Layer-Level Concepts in Terms of Channel Parameters
We found no strong clustering structure in the data. When HDBSCAN parameters were tuned to minimize unclustered samples, nearly all data fell into a single large cluster (Figure 6; bottom of Figure 7). Internal clustering quality metrics—DBCV and persistence scores—were consistently near zero (Appendix H), indicating weak and unstable groupings. Most samples formed one incoherent cluster, with remaining clusters small and likely spurious.
In contrast, 2D UMAP embeddings revealed a continuous manifold shaped by SNR (Figure 7, top). Under the baseline, data followed a single SNR-driven path. The proposed method produced two distinct paths that converge at low SNR, forming a unified structure. SNR clearly dominates the activation space, with the channel model introducing variation at high SNR and minor deviations from the main manifold among high delay spread samples. The dominance of SNR—and the limited influence of UE speed and delay spread—aligns with our replication using those parameters (Appendix G) with methodology of Tuononen et al. (2025).
Traditional receivers adapt continuously to SNR variations in filtering, demodulation, and decoding, with no discrete thresholds separating processing regimes (Tse and Viswanath (2005); Goldsmith (2005)). The continuous SNR-driven manifold observed in the activation space reflects this, suggesting the model has internalized the gradual nature of signal degradation. This alignment with classical design principles supports the interpretability and physical plausibility of the learned representations.
4.4 Evaluating Clustering Validity by using Out-of-Distribution Detection Task
We aimed to maximize the number of clustered samples, but a high clustering rate is only meaningful when paired with high clustering quality Campello et al. (2013). As this is exploratory analysis, ground truth labels are unavailable, and internal metrics like DBCV or persistence reflect algorithmic performance, not practical utility. Cross-method comparisons are limited by differing preprocessing and metrics. As noted by Jain (2010), clustering utility is domain-specific and application-dependent.
We applied the clustering structure to an Out-of-Distribution (OOD) detection task, assuming OOD samples would fall outside the clustering. Results in Figure 8 support this: the baseline clustered 15.6% of OOD samples; the proposed method, nearly none (2 samples). On In-Distribution (ID) test data, the proposed method had a lower clustering rate (79.0%) than the baseline (91.8%), suggesting better uncertainty handling and less rigid generalization. In other words, it avoids overconfidently clustering unfamiliar patterns, leading to stronger separation between ID and OOD samples. This serves as task-specific validation of clustering quality. The detection task is also practically motivated: results from over-the-air validation (Luostari et al. (2025)) suggest that general-purpose receivers must be able to detect and adapt to distribution mismatch. Further evaluation across diverse scenarios is needed to confirm broader effectiveness.
Figure 7: UMAP embeddings of DeepRx model activations from layer B1-PRE using the baseline (left) and proposed (right) methodology, colored by channel parameters and clustering partitions with HDBSCAN clustering (EoM, minimum cluster size 5). UMAP is used solely for visualization.
Figure 8: UMAP visualization of DeepRx model activations from layer B1-PRE, clustered using HDBSCAN (EoM, minimum cluster size 5). UMAP is used solely for visualization. Circles denote training samples used for clustering; crosses are test samples assigned via HDBSCAN’s predict. Left: baseline methodology; right: proposed methodology. Top: OOD test data (CDL-A, 1000 ns delay spread); bottom: ID test data (CDL channels within DeepRx training range). Visualization illustrates the extent to which ID and OOD test samples fall within or outside the learned clusters.
5 Related Work
Interpreting neural networks has led to methods for analyzing learned representations. This section reviews concept discovery approaches, including neuron-level and human-interpretable methods, clustering of distributed activation patterns, and clustering evaluation.
5.1 Concept Discovery in Neural Networks
Approaches to concept discovery often focus on either individual neurons or broader, human-interpretable concepts but may overlook layer-wide patterns emerging from distributed interactions.
Neuron-centric methods analyze individual neurons to identify features that excite specific units. Techniques include saliency maps (Simonyan et al. (2014)), which highlight input features relevant to activations; neuron maximization (Erhan et al. (2009)), which generates inputs that strongly activate specific neurons; activation histograms (Yosinski et al. (2015)), which summarize neuron behaviors; and input perturbation testing (Fong and Vedaldi (2017)), which evaluates the effects of input alterations on activations or outputs. Grad-CAM (Selvaraju et al. (2017)) produces class-specific visualizations by localizing important regions but focuses on localized activations. Network Dissection (Bau et al. (2017)) aligns neuron behaviors with predefined semantic categories (e.g., textures, objects) but fails to account for distributed patterns across layers.
Concept-based methods take a higher-level perspective by linking input features to human-interpretable concepts (Yeh et al. (2022)). Testing with Concept Activation Vectors (TCAV) (Kim et al. (2018)) measures how user-defined concepts (e.g., stripes, wheels) influence predictions, while automated approaches (Ghorbani et al. (2019)) reduce the need for manual definition. These methods have been extended to tabular data by Pendyala and Choi (2022), but they rely on predefined or discovered concept sets and may miss layer-wide patterns or concept interactions.
Modern deep networks often learn distributed and poly-semantic features, where neurons respond to multiple unrelated stimuli or encode several sub-features (Olah et al. (2017, 2020); Bau et al. (2017, 2020)). This highlights the limitations of focusing on individual neurons or predefined concepts, as individual units may learn similar or no new concepts, or interact with each other to learn a concept (Kim et al. (2018); Ghorbani et al. (2019); Bäuerle et al. (2022)).
Neural Activation Pattern (NAP) methodology (Bäuerle et al. (2022)) addresses these limitations by identifying groups of inputs with similar activation profiles across a layer, capturing layer-level concepts without retraining the model. It clusters activation patterns to reveal distributed features and visualizes them via the NAP Magnifying Glass, which presents representative images, activation statistics, and metadata filtering for comparison across layers and models, as demonstrated in a visual object recognition task by Bäuerle et al. (2022). NAP complements other interpretability methods and supports use cases such as TCAV-based testing or ante-hoc training, where interpretability is integrated during model development (Gupta et al. (2023)). By leveraging predefined concepts, NAP has the potential to enhance both interpretability and model performance.
5.2 Clustering Methods for Concept Discovery
Clustering is an unsupervised learning task that partitions data such that instances within the same cluster are more similar to each other than to those in different clusters (Jain and Dubes (1988)). Selecting an effective clustering method is essential for concept discovery, as it plays a critical role in uncovering patterns within high-dimensional activation spaces.
Hierarchical clustering creates a tree-like partition hierarchy. Methods can be categorized into agglomerative (bottom-up; e.g., BIRCH Zhang et al. (1996)) and divisive (top-down; e.g., PNN Virmajoki (2004)) approaches. Agglomerative methods iteratively merge clusters based on linkage criteria such as single linkage (shortest distance), complete linkage (farthest distance), average linkage (mean distance), or Ward’s method (minimizing variance increase). Divisive methods recursively split clusters and can be monothetic (splitting by a single attribute) or polythetic (multiple attributes).
Partitional clustering divides data into non-overlapping clusters and methods can be categorized based on the framework proposed by Tuononen (2006). This categorization includes optimization-based methods (e.g., K-Means Steinhaus (1956); MacQueen (1967); Lloyd (1982)) that minimize an objective function such as intra-cluster variance; graph-theoretic methods (e.g., spectral clustering Ng et al. (2001)) that partition similarity graphs; density-based methods (e.g., DBSCAN Ester et al. (1996)) that identify dense regions; grid-based methods, which discretize the feature space; and model-based methods (e.g., Gaussian Mixture Models Dempster et al. (1977), Self-Organizing Maps Kohonen (1990)) that fit statistical or neural network models.
HDBSCAN (Campello et al. (2013)) combines density-based and hierarchical clustering by using a minimum spanning tree for single-linkage agglomeration, effectively handling noise and irregular boundaries—ideal for high-dimensional data like neural activations with unknown cluster counts. Its accelerated variant, HDBSCAN* (McInnes and Healy (2017)), improves efficiency through optimized neighbor searches and pruning, making it well-suited for large-scale clustering tasks. The NAP framework (Bäuerle et al. (2022)) adopts this approach to reveal complex activation structures.
5.3 Evaluation of Cluster Validity
Evaluating cluster validity is essential in clustering-based concept discovery, as it determines whether the identified groupings reflect meaningful structure in the data. While many validation methods exist (Hassan et al. (2024)), choosing the right one is nontrivial: the utility of a clustering often depends highly on the specific task and domain. As Jain (2010) noted, clustering is in the eye of the beholder.
Cluster validation criteria fall into three categories (Jain and Dubes (1988)): external (compare to ground truth labels, e.g., Rand index Rand (1971), Jaccard index Jaccard (1901)); internal (assess intrinsic structure, e.g., Silhouette score Rousseeuw (1987), Davies–Bouldin score Davies and Bouldin (1979)); and relative (compare multiple clustering results, e.g. Halkidi et al. (2001)). Key indicators of cluster quality are (Jain and Dubes (1988)) compactness (minimizing distances within clusters) and separation (maximizing distances between clusters). For density-based clustering algorithms, internal validation criteria should favor high-density clusters with low-density separation. Two widely used internal indices for density-based clustering are CDbw (Halkidi and Vazirgiannis (2008)), which uses boundary points to assess structure, and DBCV (Moulavi et al. (2014)), preferred for its more accurate density estimation and boundary-free formulation.
6 Conclusions
We evaluated the Neural Activation Pattern (NAP) methodology for discovering layer-level concepts by clustering full-layer activation distributions. In both visual object recognition and radio receiver experiments, the baseline methodology with recommended HDBSCAN parameters resulted in a majority of samples remaining unclustered, with rates exceeding 90% in some cases as data volume increased. The use of Excess of Mass (EoM) as a cluster selection criterion effectively reduced this issue, while our proposed improvements—spanning normalization, distribution estimation, and distance metrics—led to more stable clustering performance across configurations.
Although distinct layer-level concepts were not identified, a continuous manifold shaped by Signal-to-Noise Ratio (SNR) consistently emerged in 2D UMAP embeddings. This structure aligns with classical receiver behavior, suggesting physical plausibility and highlighting SNR as a dominant learned factor in the receiver model’s internal representation. Our proposed methodology also improved separation between In-Distribution (ID) and Out-of-Distribution (OOD) samples, indicating stronger generalization and supporting OOD detection as both a practical application—motivated by real-world deployment needs—and an indirect measure of clustering quality. Limitations include the lack of analysis of variability due to neural network initialization and training stability, as well as the limited evaluation of OOD detection, which was tested in only a single scenario.
These findings suggest several directions for future work. The SNR-shaped manifold could support improved OOD detection through SNR estimation in receiver systems and nearest-neighbor reasoning in activation space, following runtime monitoring approaches such as Cheng et al. (2019). Additionally, since single-layer analysis did not reveal concepts for other channel parameters (e.g., delay spread), exploring concept emergence across multiple successive layers could uncover more complex structures—potentially guided by the Information Bottleneck framework by Shwartz-Ziv (2022).
Overall, our results highlight the importance of robust clustering preprocessing and parameter selection for concept discovery, with practical implications for interpretability, model troubleshooting, and runtime monitoring in deep neural networks—particularly in safety-critical, high-stakes settings such as 6G physical-layer wireless communication.
Acknowledgments and Disclosure of Funding
Ville Hautamäki was partially supported by Jane and Aatos Erkko Foundation. The authors would like to thank the creators of DeepRx and NAP framework for their foundational work. The authors would also like to thank Mikko Honkala and Heikki Penttinen for their contributions.
Supplementary Material
[…]
Appendix A Baseline Neural Activation Pattern (NAP) Methodology
The Neural Activation Pattern (NAP) methodology (Bäuerle et al. (2022)) addresses the limitations of focusing on individual neurons or predefined concepts—as described in Section 5—by identifying groups of inputs with similar activation profiles across a neural network layer. As a model-specific method, it relies on inner model details and provides post-hoc layer-level concept-based interpretability. The key steps in the NAP methodology are as follows:
  • Data Extraction: Given the layer of interest $l$ for neural network model $n$ and the set of inputs $X=\{x_{i}\ \mid i\in[1,\ldots,N]\}$ to the model, the activation output $A^{l}=\{a^{l}_{i}\mid i\in[1,\ldots,N]\}$ of the sub-network $n^{l}$ is extracted, denoted as $A^{l}=n^{l}(X)$. For convolutional layers, each $a^{l}_{i}$ is a collection of matrices (one per convolutional channel). There are no formal requirements on the input $X$ beyond those set by the neural network model. It can be the training set, test set, subsets thereof, or entirely separate data sets.
  • Location Disentanglement: For convolutional layers, spatial information in activation values can act as a confounding factor (Bäuerle et al. (2022)). To address this, activations of each convolutional channel are aggregated using statistical measurements, either maximum value (highest activation), minimum and maximum values (range of activations), mean value (average amount of activations), or mean and standard deviation (average amount and spread of activations). The disentangled activations are denoted as $\tilde{A}^{l}=\{\tilde{a}^{l}_{i}\mid i\in[1,\ldots,N]\}$. Aggregating activations removes spatial information and reduces computational cost, making the method scalable for larger models.
  • Normalization: The aggregated activation outputs are normalized by their maximum absolute activation across all inputs $X$. This max-abs scaling is given by Equation (8), where $\tilde{a}^{l}_{i,u}$ represents the disentangled activations for layer $l$ and convolutional channel $u$ with input data $x_{i}$. Channels that can only activate positively are normalized between 0 and 1, while channels with negative activations are normalized between -1 and 1.
    $\hat{a}^{l}_{i,u}=\frac{\tilde{a}^{l}_{i,u}}{\max(\left|\tilde{a}^{l}_{u}(X)\right|)}$ (8)
  • Clustering Analysis: The aggregated and normalized activation outputs $\hat{A}^{l}$ are clustered, forming a clustering $C^{l}=\{c^{l}_{j}\mid j\in[1,\ldots,M]\}$ with $M$ clusters, using the HDBSCAN* (McInnes and Healy (2017)). The clustering process employs the Euclidean distance metric, selects clusters from the leaves of the cluster tree, and uses the minimum cluster size of five. Each cluster $c^{l}_{j}\in C^{l}$ groups together a set of inputs $x_{i}\in X$ which produce similar activations for layer $l$. Clusters are mutually exclusive but not collectively exhaustive, meaning that no input $x_{i}$ can belong to more than one cluster, but some inputs $x_{i}$ may not belong to any cluster (i.e., they are classified as noise).
  • Post-Clustering Analysis: HDBSCAN* provides information about the persistence of the clusters. Persistence score reflects the stability or strength of a cluster; higher persistence meaning cluster being more well-defined and robust. Clusters are sorted based on their persistence scores, which help identify the most stable and reliable clusters. Clusters are considered to be just "raw mathematical groupings," while NAPs in general are clusters that reflect features learned by the network. Concepts are interpretable NAPs that align with human understanding.
Appendix B Neural Activation Pattern (NAP) Methodology Implementation Details
Although Bäuerle et al. (2022) mention the availability of their NAP methodology as a Python package, we were unable to locate a public implementation. Consequently, we reimplemented both their baseline approach (Appendix A) and our proposed enhancements (Section 2) using various Python libraries from which the most essential can be found from Listing 5.
In our methodology, the density is estimated using a Kernel Density Estimator (KDE) with a Gaussian kernel $K$. To obtain a compact yet informative representation, the estimated probability density function $\hat{f}^{l}_{i,u}$ (Equation 2) is sampled at $R=10$ evenly spaced Z-scores (Equation 5) over the interval $[z_{\text{min}},z_{\text{max}}]=[-2.576,2.576]$, which contains $P=99\%$ of the mass of a standard normal distribution (Equation 4). This one-point-per-decile grid (step $\Delta\approx 0.515$) yields a ten-dimensional feature vector per unit—an order-of-magnitude increase in information over a single-mean summary in baseline methodology, while remaining compact enough for downstream modeling or tabular presentation.
Appendix C Visual Object Recognition Model (ResNet50 with Pretrained Weights) and Data (ImageNet ILSVRC2012)
We employed a ResNet-50 convolutional neural network pre-trained on the ImageNet-1K dataset—namely IMAGENET1K_V2 weights from TorchVision (Marcel and Rodriguez (2010))—to perform visual object classification. To align with the experimental protocol of Bäuerle et al. (2022), we used a randomly selected 1% subset (12,832 images) of the ImageNet ILSVRC2012 training set (Russakovsky et al. (2015)). As part of the setup—and following their methodology—we extracted feature activations from the Conv4 and Conv5 stages of the network—corresponding to the layer3 and layer4 modules in TorchVision’s implementation—which produce intermediate activations of size [1024 × 14 × 14] and [2048 × 7 × 7], respectively, when processing 224 × 224 resolution images.
Appendix D Radio Receiver Model and Data (DeepRx Trained with Simulated Data)
DeepRx (Honkala et al. (2021)) is a fully convolutional deep learning model that jointly performs the tasks of channel estimation, equalization, and demapping, as illustrated in Figure 9. Moreover, the model is designed to be compliant with 5G NR specifications, capable of estimating the log-likelihood ratio (LLR), or the uncertainty, of each received bit so that it can be fed into an LDPC decoder. The model supports different demodulation reference signal (DMRS) or pilot configurations included in the 5G NR specifications (Honkala et al. (2021)).
Figure 9: Illustration of the PUSCH simulator (adapted from Honkala et al. (2021)), where DeepRx is used to replace channel estimation, equalization, and demapping blocks.
D.1 Data Simulation
The data are randomly generated using a 5G-compliant physical uplink shared channel (PUSCH) simulator, powered by NVIDIA Sionna (Hoydis et al. (2023)), and configured with the parameters listed in Table 1. The simulation process is presented in Figure 9 and consists of three main stages:
Table 1: Simulation parameters used for DeepRx training, as well as NAP training and testing. In-Distribution (ID) testing for NAP uses parameters similar to those used during NAP training. Out-of-Distribution (OOD) testing uses a fixed 1000 ns delay spread that exceeds the range seen during DeepRx training but remains within realistic bounds—specifically, corresponding to a very long delay spread as defined in Table 7.7.3-1 of ETSI (2020).
Parameter DeepRx training NAP training NAP testing (OOD)
and testing (ID)
Uniformly Distributed Channel Parameters
Signal-to-Noise Ratio (SNR) -10 dB to 25 dB -10 dB to 25 dB -10 dB to 25 dB
Doppler Shift 0 Hz to 330 Hz 0 Hz to 300 Hz 0 Hz to 300 Hz
Delay Spread 10 ns to 300 ns 10 ns to 300 ns Fixed 1000 ns
Channel Models UMi, UMa CDL-A, E for train CDL-A
All CDL for test
Other Parameters
Number of PRBs 16 (192 subcarriers)
Subcarrier Spacing 30 kHz
TTI Length 14 OFDM symbols / 0.5 ms
Modulation Scheme 16-QAM
RX / TX Antennas 2 / 1
Pilot Symbols per Slot 1 or 2 OFDM symbols
Pilot Symbol Distribution Uniformly distributed, fixed locations
  • Transmitter:
    1. 1.
      Generate random bits.
    2. 2.
      Encode the bits using a 5G-compliant LDPC encoder.
    3. 3.
      Map the encoded bits to symbols and distribute them across available physical resource blocks (PRBs) within the transmission time interval (TTI).
    4. 4.
      Insert pilot symbols, or DMRS, into the specified subcarriers.
    5. 5.
      Convert the PRBs to an OFDM waveform with 14 OFDM symbols per TTI by performing an inverse Fourier transform (IFFT).
    6. 6.
      Add a CP to the beginning of each OFDM symbol to mitigate inter-symbol interference.
  • Channel:
    1. 1.
      Pass the waveform through a channel model.
    2. 2.
      Randomly select the maximum Doppler shift and root mean square (RMS) delay spread for each channel realization.
    3. 3.
      Add noise to the signal.
  • Receiver:
    1. 1.
      Receive the signal.
    2. 2.
      Remove the CP.
    3. 3.
      Perform FFT on each OFDM symbol to convert them into the frequency domain.
    4. 4.
      Feed the frequency-domain signal of 14 OFDM symbols into the DeepRx model.
    5. 5.
      Decode the LLR output using a 5G-compliant LDPC decoder.
D.2 Model input
To maximize the network’s utilization of the input data, the model is provided with the received signal represented in the frequency domain over the entire transmission time interval (TTI). Denoting $S=14$ as the number of OFDM symbols, $F=192$ as the number of subcarriers, and $N_{r}=2$ as the number of receive antennas, the input consists of the following three arrays:
  1. 1.
    Received signal after the FFT (data and received pilot symbols), denoted by $\mathbf{Y}\in\mathbb{C}^{S\times F\times N_{r}}$.
  2. 2.
    Pilot reference symbols within the received signal Y in both frequency and time, denoted by $\mathbf{X_{p}}\in\mathbb{C}^{S\times F\times 1}$.
  3. 3.
    Raw channel estimate which is the element-wise product of the received signal $\mathbf{Y}$ and complex conjugate of the transmitted pilot symbols $\mathbf{X_{p}}$, denoted by $\mathbf{H_{r}}\in\mathbb{C}^{S\times F\times N_{r}}$.
D.3 Model architecture
The architecture of the DeepRx model, based on the work by Honkala et al. (2021) and Korpi et al. (2021), is shown in Table 2. It consists of multiple pre-activation ResNet blocks—an extension of the ResNet architecture by He et al. (2016)—and several convolutional layers. The pre-activation ResNet block, illustrated in Figure 10, is composed of two consecutive groups, each consisting of a batch normalization layer, a Rectified Linear Unit (ReLU) layer, and a depthwise separable convolution layer (Chollet (2017)). In each depthwise separable convolution block, the depth multiplier value is set to 1, meaning the number of output channels in the depthwise convolution step is the same as the input. The data flow direction is depicted in Figure 11. Certain architectural details have been omitted as they are not essential to the core contributions of our work.
The layers of interest in the paper are the ResNet blocks B1 through OUT, as illustrated in Figure 11. We focus particularly on the activations after their ReLU layers. Since each ResNet block contains two ReLU layers, we denote them as PRE and POST accordingly. For example, when referring to the second ReLU layer of ResNet block B3, we use the notation B3-POST.
Table 2: The DeepRx CNN ResNet architecture, with all convolutional layers having a filter size of (3, 3). The ResNet block is illustrated in Figure 10, and the overall data flow is depicted in Figure 11.
Layer Type Number of filters Dilation size (S, F) Output Shape
Input 1 $\mathbf{X_{p}\in\mathbb{C}}$ TX Pilot $(14,192,1)$
Input 2 $\mathbf{H_{r}\in\mathbb{C}}$ Raw channel estimate $(14,192,2)$
Input 3 $\mathbf{Y\in\mathbb{C}}$ RX data $(14,192,2)$
Pre-DeepRx Neural preprocessing $(14,192,20)$
Convolutive layer A0 2D convolution 40 (1,1) $(14,192,40)$
ResNet block A1 Depthwise separable convolution 40 (1,1) $(14,192,40)$
ResNet block A2 Depthwise separable convolution 40 (2,3) $(14,192,40)$
ResNet block A3 Depthwise separable convolution 32 (3,6) $(14,192,32)$
Convolutive layer B0 2D convolution 64 (1,1) $(14,192,64)$
ResNet block B1 Depthwise separable convolution 64 (1,1) $(14,192,64)$
ResNet block B2 Depthwise separable convolution 64 (2,4) $(14,192,64)$
ResNet block B3 Depthwise separable convolution 32 (3,8) $(14,192,32)$
ResNet block B4 Depthwise separable convolution 32 (2,4) $(14,192,32)$
ResNet block B5 Depthwise separable convolution 32 (1,1) $(14,192,32)$
ResNet block OUT Depthwise separable convolution 8 (1,1) $(14,192,8)$
LLR Output $\mathbf{L}$, see Section D.4 Output $(14,192,8)$
Figure 10: DeepRx’s pre-activation ResNet block (adapted from Honkala et al. (2021)).
Figure 11: Illustration of the data flow within the DeepRx model architecture.
D.4 Model Output and Loss Function
Since DeepRx (Honkala et al. (2021)) is designed to support multiple Quadrature Amplitude Modulation (QAM) schemes in 5G, including QPSK, 16-QAM, 64-QAM, and 256-QAM, which are hierarchically related (see Figure 13).
Figure 12: Modulation schemes used in 5G NR, including BPSK, Pi/2 BPSK, QPSK, 16-QAM, 64-QAM, and 256-QAM (redrawn from Zaidi et al. (2018)).
Figure 13: Performance comparison of the DeepRx model against baseline approaches across different SINR levels.
The model output is designed so that the same output bits, or log-likelihood ratios (LLRs), correspond to the same part of the constellation space, regardless of the modulation scheme used. For example, for an individual symbol, the first two output bits describe the complex quadrant of the symbol, called the 1st-order point; the next two bits define the quadrant within the previously identified quadrant, called the 2nd-order point. The process continues until all the bits are utilized. The visualization of this hierarchical process is provided in Figure 15.
Figure 14: The relationship between 16-QAM and QPSK constellation points (redrawn from Honkala et al. (2021)). When QPSK is used, only 2 bits $b_{1}$ and $b_{2}$ are used to define the complex quadrant, but if 16-QAM is used, besides the first two bits, $b_{3}$ and $b_{4}$ are also used to define the quadrant within the quadrant.
Figure 15: Masking bits based on the modulation, including both QPSK and 16QAM, where masking is selected based on the modulation. The output is masked accordingly so that unnecessary parts of the output array are not considered (redrawn from Honkala et al. (2021)).
In practice, the output length for each symbol is set according to the highest supported modulation order. Depending on the QAM scheme, the output is masked accordingly so that unnecessary parts of the output array are not considered. For example, if the highest modulation order is 256-QAM, each symbol in the model output is represented as an array with 8 elements for 8 bits. If a lower modulation order like 16-QAM, which requires 4 binary values, is used, the last 4 bits, or elements, are masked. The masking process is illustrated in Figure 15.
Since the bit prediction can be framed as a binary classification problem, and the network output represents the LLR of the transmitted bits, the binary cross-entropy (CE) loss is used as the loss function for the DeepRx model:
$\text{CE}(\bm{\theta})=-\frac{1}{\#DB}\sum_{(i,j)\in\mathcal{D}}\sum_{l=0}^{B-1}\left(b_{ijl}\log(\hat{b}_{ijl})+(1-b_{ijl})\log(1-\hat{b}_{ijl})\right),$ (9)
where #$\#\mathcal{D}$ is the number of resource elements carrying data; $B$ is the number of bits in the constellation used ($B=4$ for 16-QAM); $b_{ijl}$ is the label bit; and $\hat{b}_{ijl}$ is the probability that bit $b_{ijl}$ equals one, calculated using the sigmoid function:
$\hat{b}_{ijl}=\text{sigmoid}\left(L_{ijl}\right)=\frac{1}{1+e^{-L_{ijl}}}.$ (10)
D.5 Model’s Predictive Performance
The DeepRx model demonstrates competitive performance in terms of uncoded Bit Error Rate (BER) across varying Signal-to-Interference-plus-Noise-Ratio (SINR) levels, as seen in Figure 13. It closely matches the baseline performance with a known channel, particularly at low SINR values. Notably, it outperforms the baseline with one pilot in higher SINR scenarios, which aligns with the findings reported in Honkala et al. (2021).
Appendix E Supplementary Analysis of Dataset Size Impact on Visual Object Recognition Using ResNet50 with Pretrained Weights and ImageNet Data
This appendix explores the impact of dataset size on visual object recognition using the ResNet50 architecture with pretrained weights and activations from the Conv4 and Conv5 stages of the network, as described in Appendix C. Figures 17 and 17 illustrate how varying the dataset size influences the number of clusters and the mean persistence score across different layers and methodologies.
Figure 16: Number of clusters per layer and methodology as function of sample size with minimum cluster size 5 using ImageNet subset with ResNet-50 model.
Figure 17: Mean persistence score per layer and methodology as function of sample size with minimum cluster size 5 using ImageNet subset with ResNet-50 model.
Appendix F Supplementary Analysis of Dataset Size Impact on Radio Receiver Using DeepRx with Simulated Data
This appendix explores the impact of dataset size on radio receiver using the DeepRx architecture and activations from various stages of the network, as described in Appendix D. Figures 19 and 19 illustrate how varying the dataset size influences the number of clusters and the mean persistence score across different layers and methodologies.
Figure 18: Number of clusters per layer and methodology as function of sample size with minimum cluster size 5 using link-level simulations with DeepRx model.
Figure 19: Mean persistence score per layer and methodology as function of sample size with minimum cluster size 5 using link-level simulations with DeepRx model.
Appendix G Experiments on Delay Spread and UE Speed Using the NUCPI Methodology
In this appendix, we present our replication of the experiments conducted by Tuononen et al. (2025), employing delay spread (see Appendix G.2) and user equipment (UE) speed (see Appendix G.3) as target parameters. This approach diverges from the original study, which utilized Signal-to-Noise Ratio (SNR) as the channel parameter of interest. Notably, Tuononen et al. (2025) suggested that their method could be applicable to other channel parameters, which we explore here.
G.1 NUCPI Methodology and Its Potential
We refer to the methodology proposed in Tuononen et al. (2025) as Neural Unit Channel Parameter Identification (NUCPI), wherein an explainer model is trained to predict a given channel parameter based on intermediate activations from a performer model—in our case, the DeepRx model (see Appendix D). We interpret the predictive performance of the explainer model—measured using Mean Squared Error (MSE), as suggested by Tuononen et al. (2025)—as an indicator of how much information specific internal components (e.g., layers or channels) of the performer contain about the given channel parameter. This enables insights into the model’s internal behavior.
These insights could potentially be used to identify redundant or over-specialized components, guide model pruning Han et al. (2016), or adapt architectures to better handle specific channel conditions Cortes et al. (2017). While these applications are not experimentally explored in this work, the interpretability provided by NUCPI opens promising directions for future research in model compression and adaptive architecture design.
G.2 NUCPI Analysis of Delay Spread
We performed NUCPI analysis of delay spread using both layerwise and channelwise (for layer B1-PRE) perspectives. As shown in the layerwise (Figure 20) and channelwise (Figure 21) test losses of the explainer models, predictive performance is consistently poor across all units. Upon closer examination, it is evident that the explainer models are overfitting: while training loss decreases steadily, evaluation loss plateaus or increases, indicating limited generalization to unseen data (Goodfellow et al. (2016)). This is illustrated in training curves for both the layerwise (B1-PRE in Figure 23; OUT-POST in Figure 23) and channelwise (channels 7 and 12 on B1-PRE, in Figures 25 and 25, respectively) explainer models.
The estimates produced by the explainer models (see Figure 28 for layerwise, and Figures 30 and 29 for channelwise) show that the models tend to predict values near the mean of the delay spread distribution. This pattern occurs both in layerwise and channelwise models, suggesting that the models are effectively ignoring input variations and defaulting to the values around the global average. We conclude that the activations of the DeepRx model likely do not contain meaningful information about delay spread, at least not in a directly predictive form.
From a physical layer perspective, this is plausible given our simulation parameters (see Table 1). The bandwidth is $B=192\times 30\text{ kHz}=5.76\text{ MHz}$, yielding a time resolution of $\Delta\tau=1/B\approx 174\text{ ns}$, and a quantization uncertainty of approximately $\sigma\approx\Delta\tau/\sqrt{12}\approx\pm 50\text{ ns}$ (standard deviation of a uniform distribution (Earl and Nicholson, 2021, Appendix 15)). As a result, multipath components closer than 174 ns apart (e.g., a 10 ns path) are indistinguishable, and even larger separations are affected by quantization noise. This limitation likely contributes to the lack of delay spread information in DeepRx’s activations. Interestingly, DeepRx still performs well under delay spread conditions, suggesting it may compensate for this effect without needing to explicitly quantify the delay spread (e.g., as RMS).
G.3 NUCPI Analysis of User Equipment (UE) Speed
We also applied NUCPI analysis to UE speed, evaluating layerwise explainer performance for two layers. As with delay spread (see Appendix G.2), the explainer models exhibit consistent overfitting: training losses decrease, but evaluation losses do not improve, as shown in Figures 27 and 27. Based on this behavior, we conclude that DeepRx’s internal activations likely do not encode UE speed in a directly predictive way.
In traditional systems, UE speed—or Doppler spread more technically—is estimated by aggregating over tens of slots (Yucek et al. (2005); Zhao et al. (2009)). In contrast, DeepRx operates on a single slot without a memory of past slots. Despite this, DeepRx can still handle frequency offset effects caused by user movement (and other factors). A similar explanation as with delay spread may apply: the model may not need to explicitly estimate frequency offset or Doppler spread to effectively compensate for them.
Figure 20: Layerwise test losses for the explainer model on delay spread, showing poor predictive performance across all layers.
Figure 21: Channelwise test losses for the explainer model on delay spread for layer B1-PRE, showing poor predictive performance across all channels.
Figure 22: Training curve on delay spread for layer B1-PRE, illustrating overfitting in the explainer model.
Figure 23: Training curve on delay spread for layer OUT-POST, showing overfitting in the explainer model.
Figure 24: Training curve on delay spread for channel 7 on layer B1-PRE, indicating overfitting in the explainer model.
Figure 25: Training curve on delay spread for channel 12 on layer B1-PRE, illustrating overfitting in the explainer model.
Figure 26: Training curve on UE speed for layer B1-PRE, illustrating overfitting in the explainer model.
Figure 27: Training curve on UE speed for layer OUT-POST, illustrating overfitting in the explainer model.
Figure 28: Estimates of delay spread by layerwise explainer models, peaking around the mean of true delay spread parameters.
Figure 29: Estimates of delay spread by channelwise explainer models on layer OUT-POST, peaking around the mean of true delay spread parameters.
Figure 30: Estimates of delay spread by channelwise explainer models on layer B1-PRE, peaking around the mean of true delay spread parameters.
Appendix H Cluster Statistics and DBCV Scores for the Example Case
This appendix presents detailed cluster statistics and DBCV scores for the example case shown in Figure 7. The baseline method results are shown in Listings 1 and 3, while the proposed method results are in Listings 2 and 4.
Cluster #0, 7 samples
Persistence score: 0.01
Data probability: 1.00 +- 0.01
SNR: 7.60 +- 11.97 dB
UE speed: 71.76 +- 31.59 km/h
Delay spread: 111.52 +- 72.39 ns
Channel model: 0.43 +- 0.49
CDL-A (nlos): 7 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 0 samples
BER: 0.0669 +- 0.0746
SER: 0.3297 +- 0.3387
Cluster #1, 38373 samples
Persistence score: 0.12
Data probability: 0.98 +- 0.05
SNR: 7.50 +- 10.10 dB
UE speed: 62.08 +- 35.99 km/h
Delay spread: 155.57 +- 83.73 ns
Channel model: 0.50 +- 0.50
CDL-A (nlos): 18686 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 19687 samples
BER: 0.0569 +- 0.0649
SER: 0.3004 +- 0.3065
Cluster #2, 6 samples
Persistence score: 0.00
Data probability: 1.00 +- 0.00
SNR: 10.09 +- 9.24 dB
UE speed: 57.99 +- 24.47 km/h
Delay spread: 155.65 +- 89.27 ns
Channel model: 0.50 +- 0.50
CDL-A (nlos): 6 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 0 samples
BER: 0.0402 +- 0.0383
SER: 0.2617 +- 0.2329
Listing 1: Cluster statistics for baseline methodology in example case in Figure 7
Cluster #0, 14 samples
Persistence score: 0.00
Data probability: 1.00 +- 0.00
SNR: 6.52 +- 11.25 dB
UE speed: 47.24 +- 32.21 km/h
Delay spread: 144.79 +- 75.15 ns
Channel model: 0.71 +- 0.45
CDL-A (nlos): 8 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 6 samples
BER: 0.0643 +- 0.0719
SER: 0.3190 +- 0.3241
Cluster #1, 5 samples
Persistence score: 0.00
Data probability: 1.00 +- 0.00
SNR: 10.01 +- 9.96 dB
UE speed: 90.18 +- 25.18 km/h
Delay spread: 221.67 +- 58.79 ns
Channel model: 0.40 +- 0.49
CDL-A (nlos): 2 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 3 samples
BER: 0.0496 +- 0.0609
SER: 0.2647 +- 0.3241
Cluster #2, 5 samples
Persistence score: 0.00
Data probability: 1.00 +- 0.00
SNR: 4.45 +- 9.56 dB
UE speed: 63.30 +- 16.96 km/h
Delay spread: 198.46 +- 30.81 ns
Channel model: 0.60 +- 0.49
CDL-A (nlos): 3 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 2 samples
BER: 0.0745 +- 0.0674
SER: 0.3771 +- 0.3149
Cluster #3, 6 samples
Persistence score: 0.00
Data probability: 1.00 +- 0.00
SNR: 10.41 +- 10.36 dB
UE speed: 68.90 +- 43.58 km/h
Delay spread: 119.16 +- 79.94 ns
Channel model: 0.50 +- 0.50
CDL-A (nlos): 2 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 4 samples
BER: 0.0485 +- 0.0584
SER: 0.2651 +- 0.3016
Cluster #4, 10 samples
Persistence score: 0.00
Data probability: 1.00 +- 0.00
SNR: 11.19 +- 8.86 dB
UE speed: 76.13 +- 37.53 km/h
Delay spread: 157.09 +- 72.29 ns
Channel model: 0.80 +- 0.40
CDL-A (nlos): 4 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 6 samples
BER: 0.0346 +- 0.0577
SER: 0.1866 +- 0.2873
Cluster #5, 6 samples
Persistence score: 0.01
Data probability: 1.00 +- 0.00
SNR: 8.93 +- 7.36 dB
UE speed: 42.42 +- 29.41 km/h
Delay spread: 185.68 +- 97.80 ns
Channel model: 0.67 +- 0.47
CDL-A (nlos): 4 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 2 samples
BER: 0.0360 +- 0.0425
SER: 0.2275 +- 0.2441
Cluster #6, 33063 samples
Persistence score: 0.07
Data probability: 1.00 +- 0.01
SNR: 7.45 +- 10.11 dB
UE speed: 61.97 +- 36.01 km/h
Delay spread: 155.05 +- 83.93 ns
Channel model: 0.50 +- 0.50
CDL-A (nlos): 16415 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 16648 samples
BER: 0.0572 +- 0.0651
SER: 0.3021 +- 0.3070
Cluster #7, 6 samples
Persistence score: 0.00
Data probability: 1.00 +- 0.00
SNR: 4.10 +- 10.52 dB
UE speed: 55.36 +- 27.38 km/h
Delay spread: 139.33 +- 69.94 ns
Channel model: 0.50 +- 0.50
CDL-A (nlos): 2 samples
CDL-B (nlos): 0 samples
CDL-C (nlos): 0 samples
CDL-D (los): 0 samples
CDL-E (los): 4 samples
BER: 0.0837 +- 0.0712
SER: 0.4219 +- 0.3209
Listing 2: Cluster statistics for proposed methodology in example case in Figure 7
DBCV for clustering: 0.0555
DBCV for cluster #0: 0.1843
DBCV for cluster #1: 0.0578
DBCV for cluster #2: 0.2587
Listing 3: DBCV scores for baseline methodology in example case in Figure 7
DBCV for clustering: 0.0141
DBCV for cluster #0: 0.0244
DBCV for cluster #1: 0.0311
DBCV for cluster #2: 0.0752
DBCV for cluster #3: 0.0685
DBCV for cluster #4: 0.1635
DBCV for cluster #5: 0.1868
DBCV for cluster #6: 0.0170
DBCV for cluster #7: 0.0279
Listing 4: DBCV scores for proposed methodology in example case in Figure 7
Appendix I On the Illusory Advantages of UMAP Dimensionality Reduction on the OUT-POST Layer of the DeepRx Model
While UMAP is commonly employed to enhance the separability of high-dimensional data in lower-dimensional spaces, our experiments indicate that its application to OUT-POST layer activations in the DeepRx model may introduce distortions that compromise interpretability. In particular, we suspect that UMAP may “tear” the embedding space, leading to artificially induced clustering that lacks correspondence to meaningful channel parameter variations. Figures 32 and 32 illustrate these divergent clustering outcomes. Although UMAP superficially increases cluster compactness, it appears to do so at the cost of semantic coherence—highlighting the need for caution when applying nonlinear dimensionality reduction in domains where physical interpretability is critical.
This suspected tearing behavior aligns with known limitations of UMAP. The algorithm prioritizes local neighborhood preservation under assumptions of uniform data distribution and local manifold connectivity. These assumptions can result in global structural distortions and discontinuities in the embedded space (McInnes et al. (2018)). Coenen et al. (2019) provide an intuitive visualization of this effect, showing how UMAP may artificially separate continuous regions. Damrich and Hamprecht (2021) further offer a theoretical analysis, attributing these artifacts to the optimization objective and the underlying graph construction. These insights are consistent with our observations and underscore the importance of hyperparameter tuning and critical interpretation of UMAP results—particularly in contexts where maintaining topological and semantic continuity is essential.
Figure 31: Clustering result after UMAP reduction to five dimensions. Although over half the data (24k samples) is assigned to clusters, the resulting 1,176 clusters show no clear physical interpretation. Clustering was based on EoM, minimum cluster size of 5, and proposed methodology.
Figure 32: Clustering result without UMAP. A bit fewer samples (19k samples) are clustered, but the 139 resulting clusters are more interpretable (clusters on low and high SNR ends). Clustering was based on EoM, minimum cluster size of 5, and proposed methodology.
Appendix J Examples of 2D UMAP Embeddings Using Unaggregated and Unnormalized DeepRx Activations
In this appendix, we present two-dimensional UMAP embeddings of the raw (i.e., unaggregated and unnormalized) activations from the DeepRx model (Appendix D) for the B3-POST and OUT-POST layers. As discussed in the main paper, direct clustering of these activations using HDBSCAN was unsuccessful, with all samples classified as outliers—likely due to the high dimensionality and lack of preprocessing in the feature space.
Additionally, an HDBSCAN run on the B1-PRE layer, using a single parameter configuration, failed to complete after approximately two weeks (341 hours) of execution. The exact cause of termination remains unclear, but the duration and lack of output suggest a computational bottleneck or instability in the clustering process.
While the use of UMAP occasionally enabled some clustering structure to emerge, the results were highly sensitive to both the layer and the chosen parameters, and no consistent patterns were observed. However, despite these limitations, the 2D UMAP projections of these raw activations reveal a continuous low-dimensional manifold primarily structured by SNR. This trend is visible in both B3-POST and OUT-POST layers (see Figure 34 and Figure 38) and aligns with the structure previously observed in preprocessed activations using both the baseline and proposed methodologies (Figure 7).
Figure 34 suggests that, for layer B3-POST, lower UE speeds tend to concentrate near the center of the manifold, while higher speeds deviate outward. However, this observation is based solely on visual inspection and exhibits considerable noise—some higher speed samples also appear near the center. A more rigorous, quantitative analysis would be necessary to draw definitive conclusions. This trend is not observed in the OUT-POST layer (Figure 38).
Figure 40 shows that some CDL-A samples in the OUT-POST layer deviate slightly from the main manifold, though this effect is subtle and not evident in the corresponding B3-POST visualization (Figure 36).
Finally, we observe no clear visual structure related to delay spread in either B3-POST (Figure 36) or OUT-POST (Figure 40) layers.
Figure 33: 2D UMAP embedding of raw activations from the DeepRx model’s B3-POST layer, colored by Signal-to-Noise Ratio (SNR). The manifold structure suggests an organization influenced by SNR.
Figure 34: 2D UMAP embedding of raw activations from the DeepRx model’s B3-POST layer, colored by User Equipment (UE) speed. A tendency for lower speeds to concentrate near the center is observable, though with noise.
Figure 35: 2D UMAP embedding of raw activations from the DeepRx model’s B3-POST layer, colored by delay spread values. No clear structure related to delay spread is visually apparent.
Figure 36: 2D UMAP embedding of raw activations from the DeepRx model’s B3-POST layer, colored by channel model. No distinct separation between different channel models is observed.
Figure 37: 2D UMAP embedding of raw activations from the DeepRx model’s OUT-POST layer, colored by Signal-to-Noise Ratio (SNR). Similar to B3-POST, a clear SNR-driven structure is evident.
Figure 38: 2D UMAP embedding of raw activations from the DeepRx model’s OUT-POST layer, colored by User Equipment (UE) speed. No consistent pattern associated with UE speed is observed.
Figure 39: 2D UMAP embedding of raw activations from the DeepRx model’s OUT-POST layer, colored by delay spread values. Delay spread does not appear to influence the structure of the embedding.
Figure 40: 2D UMAP embedding of raw activations from the DeepRx model’s OUT-POST layer, colored by channel model class. Some CDL-A samples show minor deviation from the main manifold.
Appendix K Compute Resources and Environment
We outline the compute resources, memory and storage specifications, software stack, and execution time for our experimental setup. The execution time figures are approximate, derived from logs, rather than precise measurements from instrumentation or formal profiling tools.
K.1 Compute Environment
All experiments were conducted using the following cloud-based compute resources:
  • GPU: Multi-Instance GPU (MIG) instances, each providing 10GB of dedicated GPU memory. The performance of each instance is approximately equivalent to a lower-end dedicated GPU such as the NVIDIA T4, or roughly one-seventh (1/7th) of a high-end data center GPU (NVIDIA A100 80GB).
  • CPU: 16 virtual CPU cores. These cores are comparable to those from 3rd Generation Intel Xeon Scalable processors (Ice Lake architecture), commonly found in enterprise and data center environments.
  • Memory: 512 GB system RAM
  • Storage: 1 TB HDD
  • Software Stack: Python 3.10, PyTorch 2.6, TensorFlow 2.6, and many other dependencies. The most essential dependencies are specified in Listing 5.
h5py==3.13.0
hdbscan==0.8.40
joblib==1.4.2
matplotlib==3.10.1
numba==0.61.0
numpy==2.1.3
pandas==2.2.3
scikit-learn==1.6.1
scipy==1.15.2
sionna==0.10.0
torch==2.6.0
torchvision==0.21.0
umap-learn==0.5.7
tensorflow==2.6.4
Listing 5: Most essential Python requirements used in requirements.txt format.
K.2 Compute Resources for Reported Experimental Runs
Table 3 summarizes the estimated execution times for the overall experimental results presented in the main paper and for the detailed experimental results presented in Appendix L.
Table 3: Estimated execution times for reported experiments.
Experiment Execution Time (in hours)
(Prerequisite) Training and evaluating DeepRx model
and preprocessing activations using 1 MIG
42
NAP on Visual Object Recognition – Dataset Size 18
NAP on Visual Object Recognition – HDBSCAN Parameters 13
NAP on Radio Receiver – Dataset Size 14
NAP on Radio Receiver – HDBSCAN Parameters 782
NAP on Radio Receiver – UMAP Dimensionality Reduction 208
Estimated Total 1077
K.3 Compute Resources for Additional Experiments
In addition to the main experiments, we conducted several exploratory, debugging, and preliminary experiments to guide methodology development. These are summarized in Table 4. Estimated totals reflect rough usage and include failed and unreported experiments that were discarded, did not yield improvements, or were otherwise not included in the main paper.
Table 4: Estimated execution times for additional experimental work.
Experiment Execution Time (in hours)
Preliminary work on NAP methodology (Master’s thesis) 1030
NUCPI analysis using 1 MIG (Appendix G) 144
Debugging, sanity checks, and failed and discarded experiments 1052
Estimated Total 2226
Appendix L Detailed Results from HDBSCAN Parameter Sweeps and UMAP Dimensionality Reduction Experiments on DeepRx
This appendix presents detailed tabular results (Tables LABEL:tab:b1_pre_non_clustered_eom to LABEL:tab:out_post_mean_persistence_leaf) from experiments evaluating the impact of HDBSCAN clustering parameters, with and without UMAP-based dimensionality reduction, on various preprocessing strategies applied to DeepRx (Radio Receiver model; Appendix D) activations. The main paper and other appendices summarize and discuss key findings drawn from these tables.
Note on UMAP dimensionality. Healy and McInnes (2024) notes that dimensionality reduction can benefit clustering, particularly when reducing to more than three dimensions—typically between 4 and 10—and that the number of reduced dimensions should generally not exceed the n_neighbors parameter. Following this guidance, we experimented with 5, 10, and 15 dimensions, aligning with common UMAP practices and staying within the default n_neighbors value of 15.
Table 5: Number of Non-Clustered Samples Using Excess of Mass (EoM) Criteria on Layer B1-PRE
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 17644 17929 18081 316 18139 18745 19012 721 25273 29232 116 -
5 20789 19940 21703 6885 21484 22393 22875 1614 2333 1755 25857 -
7 21620 20606 21398 22137 39 23148 24115 11007 27218 26711 26682 -
9 15 21635 20579 24623 45 45 54 8659 28494 28440 25578 -
11 19279 21657 20333 18543 23179 682 886 8935 24884 26239 27018 -
13 21222 21798 19566 18595 583 22809 24152 37495 28434 25278 25858 -
15 0 19069 21497 18389 23280 23325 1059 37764 25829 24107 27048 -
17 19950 22167 21773 18699 15584 20836 2395 37126 25409 25950 26706 -
19 20409 13387 0 19182 13379 23188 25076 37334 28074 25563 26586 -
Table 5: Number of Non-Clustered Samples Using Excess of Mass (EoM) Criteria on Layer B1-PRE
Table 6: Number of Non-Clustered Samples Using Leaves of the Tree (Leaf) Criteria on Layer B1-PRE
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 20866 21216 21194 37283 21090 21518 21643 37151 26410 30468 31163 -
5 24705 25068 25384 39146 25297 25401 25841 38947 33312 36237 36772 -
7 26544 27046 26928 39510 26595 26854 27161 39231 35639 37526 38092 -
9 27710 27881 27831 39542 27962 27912 28103 39271 37147 37802 38252 -
11 27808 28249 27846 39594 28600 28632 29002 39410 37593 37754 38069 -
13 27596 28539 27792 39653 29050 28916 29418 39360 37837 38379 38359 -
15 27927 28257 28089 39756 29106 29356 29424 38811 37584 38264 37998 -
17 26805 28469 27977 39436 29263 29495 29437 38738 37848 38446 38388 -
19 27793 28023 27725 39838 29094 29748 29736 38802 37955 38009 38595 -
Table 6: Number of Non-Clustered Samples Using Leaves of the Tree (Leaf) Criteria on Layer B1-PRE
Table 7: Number of Clusters Using Excess of Mass (EoM) Criteria on Layer B1-PRE
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 2430 2410 2427 2 2435 2355 2409 7 2207 1735 2 -
5 1014 945 1062 8 998 992 1052 3 3 2 183 -
7 579 517 526 10 2 568 593 3 198 60 45 -
9 2 371 329 6 2 2 2 2 103 37 20 -
11 236 268 247 2 280 2 2 2 30 18 17 -
13 218 216 189 2 3 212 223 6 30 14 11 -
15 2 146 178 2 179 187 2 7 19 10 9 -
17 142 153 148 2 93 127 7 5 10 8 6 -
19 133 90 2 2 65 133 147 5 11 5 9 -
Table 7: Number of Clusters Using Excess of Mass (EoM) Criteria on Layer B1-PRE
Table 8: Number of Clusters Using Leaves of the Tree (Leaf) Criteria on Layer B1-PRE
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 2908 2913 2880 580 2916 2806 2795 585 2389 1748 1635 -
5 1331 1264 1255 91 1228 1195 1188 101 698 386 332 -
7 755 747 750 33 728 725 727 43 304 144 114 -
9 533 524 507 17 528 499 502 25 143 76 71 -
11 405 406 394 12 393 392 392 23 89 53 42 -
13 312 311 315 10 316 321 323 18 54 35 33 -
15 268 270 267 6 270 254 254 14 47 29 29 -
17 221 224 225 4 224 226 217 12 31 23 25 -
19 199 198 196 4 195 198 197 9 27 22 20 -
Table 8: Number of Clusters Using Leaves of the Tree (Leaf) Criteria on Layer B1-PRE
Table 9: Maximum Persistence Score Using Excess of Mass (EoM) Criteria on Layer B1-PRE
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 0.4340 0.4000 0.4743 0.0262 0.5282 0.2861 0.3955 0.0248 0.2429 0.1977 0.0169 -
5 0.4025 0.3544 0.5125 0.0700 0.3031 0.3219 0.3201 0.1231 0.0260 0.0083 0.0594 -
7 0.4602 0.3351 0.3450 0.0497 0.0420 0.4176 0.3319 0.0724 0.0731 0.0575 0.1317 -
9 0.0886 0.3150 0.3785 0.0652 0.0399 0.0268 0.0285 0.1440 0.0589 0.0684 0.0704 -
11 0.3248 0.2415 0.2782 0.0840 0.4218 0.0749 0.0795 0.1425 0.0696 0.0979 0.0777 -
13 0.2966 0.4422 0.2996 0.0838 0.1535 0.2098 0.3442 0.0327 0.0776 0.1082 0.1061 -
15 0.0774 0.2861 0.3366 0.1130 0.3752 0.2851 0.0237 0.0329 0.0412 0.0573 0.0956 -
17 0.3444 0.2658 0.2726 0.1147 0.2565 0.2665 0.1208 0.0416 0.0672 0.1230 0.1145 -
19 0.2966 0.3065 0.2321 0.1138 0.1782 0.2972 0.1603 0.0434 0.0770 0.1284 0.1105 -
Table 9: Maximum Persistence Score Using Excess of Mass (EoM) Criteria on Layer B1-PRE
Table 10: Maximum Persistence Score Using Leaves of the Tree (Leaf) Criteria on Layer B1-PRE
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 0.3981 0.4888 0.5531 0.0682 0.5845 0.4235 0.3600 0.0813 0.1704 0.1631 0.1573 -
5 0.3077 0.3574 0.4515 0.0259 0.3791 0.3785 0.2945 0.0338 0.1612 0.0938 0.0947 -
7 0.3310 0.4587 0.3795 0.0197 0.4222 0.4649 0.3266 0.0343 0.0686 0.1327 0.1228 -
9 0.2627 0.3022 0.4791 0.0172 0.4005 0.4033 0.3403 0.0278 0.0528 0.1380 0.1125 -
11 0.2946 0.2566 0.3372 0.0169 0.3604 0.3433 0.3331 0.0245 0.1215 0.1276 0.1015 -
13 0.2879 0.2965 0.2051 0.0148 0.3007 0.3780 0.3391 0.0271 0.1202 0.1518 0.1443 -
15 0.2872 0.3133 0.3651 0.0157 0.3589 0.3432 0.3273 0.0477 0.1374 0.1206 0.1482 -
17 0.3076 0.2575 0.2490 0.0198 0.2248 0.3208 0.2814 0.0466 0.1249 0.1316 0.1211 -
19 0.2877 0.2651 0.2787 0.0178 0.2997 0.3225 0.3150 0.0453 0.1283 0.1266 0.1301 -
Table 10: Maximum Persistence Score Using Leaves of the Tree (Leaf) Criteria on Layer B1-PRE
Table 11: Mean Persistence Score Using Excess of Mass (EoM) Criteria on Layer B1-PRE
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 0.0304 0.0159 0.0181 0.0150 0.0266 0.0207 0.0284 0.0075 0.0255 0.0162 0.0109 -
5 0.0278 0.0395 0.0302 0.0117 0.0170 0.0316 0.0272 0.0450 0.0196 0.0042 0.0066 -
7 0.0459 0.0439 0.0293 0.0126 0.0295 0.0299 0.0362 0.0272 0.0104 0.0064 0.0090 -
9 0.0587 0.0394 0.0583 0.0231 0.0336 0.0217 0.0270 0.0723 0.0095 0.0092 0.0149 -
11 0.0469 0.0457 0.0553 0.0623 0.0339 0.0487 0.0480 0.0717 0.0099 0.0175 0.0149 -
13 0.0485 0.0583 0.0566 0.0799 0.0785 0.0422 0.0432 0.0151 0.0115 0.0194 0.0231 -
15 0.0679 0.0504 0.0574 0.1001 0.0495 0.0461 0.0234 0.0152 0.0107 0.0156 0.0212 -
17 0.0546 0.0581 0.0495 0.1002 0.0361 0.0399 0.0499 0.0174 0.0177 0.0313 0.0472 -
19 0.0570 0.0648 0.1858 0.0987 0.0446 0.0545 0.0392 0.0173 0.0183 0.0623 0.0273 -
Table 11: Mean Persistence Score Using Excess of Mass (EoM) Criteria on Layer B1-PRE
Table 12: Mean Persistence Score Using Leaves of the Tree (Leaf) Criteria on Layer B1-PRE
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 0.0288 0.0232 0.0194 0.0066 0.0131 0.0197 0.0191 0.0130 0.0208 0.0124 0.0098 -
5 0.0211 0.0343 0.0318 0.0053 0.0194 0.0278 0.0312 0.0118 0.0127 0.0110 0.0098 -
7 0.0313 0.0232 0.0342 0.0063 0.0302 0.0319 0.0315 0.0127 0.0121 0.0115 0.0111 -
9 0.0362 0.0350 0.0435 0.0063 0.0323 0.0286 0.0360 0.0136 0.0103 0.0120 0.0146 -
11 0.0446 0.0408 0.0484 0.0078 0.0369 0.0386 0.0344 0.0124 0.0106 0.0112 0.0154 -
13 0.0486 0.0456 0.0389 0.0071 0.0386 0.0446 0.0401 0.0121 0.0132 0.0169 0.0150 -
15 0.0504 0.0517 0.0509 0.0064 0.0495 0.0384 0.0443 0.0129 0.0153 0.0167 0.0188 -
17 0.0577 0.0515 0.0523 0.0108 0.0326 0.0422 0.0422 0.0140 0.0152 0.0170 0.0179 -
19 0.0434 0.0530 0.0569 0.0057 0.0464 0.0403 0.0425 0.0155 0.0166 0.0190 0.0160 -
Table 12: Mean Persistence Score Using Leaves of the Tree (Leaf) Criteria on Layer B1-PRE
Table 13: Number of Non-Clustered Samples Using Excess of Mass (EoM) Criteria on Layer B3-POST
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 15927 16034 16316 374 15588 15834 16322 47 25576 334 27702 -
5 18428 19195 19192 1767 18720 19146 18831 548 2854 2026 1863 -
7 19760 19642 19930 6516 19820 19640 20579 3274 1437 31598 4497 -
9 17997 18212 20794 2041 19688 21573 21050 13107 31407 32173 3244 -
11 17247 20475 21649 5234 20878 21768 21658 12545 31394 6121 2855 40000
13 18758 18435 18250 18300 22034 22485 68 16282 4935 23034 9830 -
15 17528 17528 17042 18440 20985 22205 554 15883 2224 18771 25470 -
17 19425 18457 20174 18733 22207 33 57 13647 3748 4355 22949 -
19 18835 17834 18321 18994 528 344 22328 13865 5021 5165 18218 -
Table 13: Number of Non-Clustered Samples Using Excess of Mass (EoM) Criteria on Layer B3-POST
Table 14: Number of Non-Clustered Samples Using Leaves of the Tree (Leaf) Criteria on Layer B3-POST
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 19466 19598 19913 35723 19045 19111 19287 34625 27260 29115 29114 -
5 22678 23001 23122 37927 22285 22519 22443 37496 32087 32858 32974 -
7 23836 23917 23936 38584 23623 24051 24193 38278 34068 34603 34780 -
9 24127 24677 24538 38688 24561 25230 24813 38714 34778 35171 35524 -
11 24550 24888 24756 38877 25007 25391 25473 38829 35445 35650 36051 -
13 24932 24989 25448 38416 26171 26371 25891 38972 35842 35921 35969 -
15 25102 26020 26018 38432 26305 27116 26720 38767 35990 36318 35911 -
17 26083 26062 26014 38532 26659 27050 26816 38854 35841 36770 36876 -
19 26256 26626 26709 38635 27186 27424 27046 39063 36283 36619 36846 -
Table 14: Number of Non-Clustered Samples Using Leaves of the Tree (Leaf) Criteria on Layer B3-POST
Table 15: Number of Clusters Using Excess of Mass (EoM) Criteria on Layer B3-POST
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 2419 2429 2447 2 2527 2512 2587 2 1868 2 1557 -
5 1121 1125 1114 2 1192 1231 1170 2 3 2 2 -
7 689 679 689 6 759 703 753 2 3 269 2 -
9 455 478 517 2 526 570 558 5 193 160 2 -
11 354 375 406 3 391 428 416 3 102 2 2 0
13 297 294 287 5 341 326 2 4 2 28 2 -
15 239 240 246 4 263 269 7 4 2 15 34 -
17 227 210 234 3 233 2 2 3 3 2 19 -
19 202 187 183 4 4 4 183 3 2 2 14 -
Table 15: Number of Clusters Using Excess of Mass (EoM) Criteria on Layer B3-POST
Table 16: Number of Clusters Using Leaves of the Tree (Leaf) Criteria on Layer B3-POST
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 3035 2950 3000 798 3062 3021 3025 1056 2118 1852 1853 -
5 1390 1367 1383 198 1450 1436 1437 234 751 657 650 -
7 877 860 849 91 930 924 909 97 386 355 354 -
9 622 638 616 51 665 663 644 54 246 231 221 -
11 488 484 493 35 513 518 516 41 167 153 154 -
13 397 393 401 24 406 404 424 30 121 109 123 -
15 346 342 336 19 356 350 343 25 89 89 79 -
17 298 304 307 13 298 282 287 22 83 74 68 -
19 272 265 270 13 264 252 246 18 67 57 60 -
Table 16: Number of Clusters Using Leaves of the Tree (Leaf) Criteria on Layer B3-POST
Table 17: Maximum Persistence Score Using Excess of Mass (EoM) Criteria on Layer B3-POST
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 0.4874 0.5495 0.5570 0.0365 0.7108 0.6318 0.5955 0.0486 0.3142 0.0088 0.1945 -
5 0.3157 0.4265 0.2647 0.0226 0.3351 0.4021 0.3361 0.0724 0.0453 0.0113 0.0070 -
7 0.4558 0.4752 0.4858 0.0504 0.2923 0.3139 0.3357 0.0773 0.0169 0.1638 0.0053 -
9 0.4657 0.6086 0.4926 0.0589 0.2724 0.2978 0.3855 0.0238 0.2351 0.0772 0.0142 -
11 0.5540 0.5121 0.3155 0.0843 0.2550 0.2342 0.2771 0.0301 0.1005 0.0546 0.0186 N/A
13 0.3741 0.5287 0.4160 0.1104 0.3261 0.2864 0.1331 0.0509 0.0358 0.1071 0.0105 -
15 0.4748 0.4211 0.3416 0.1124 0.2413 0.2375 0.1075 0.0592 0.0314 0.0476 0.1232 -
17 0.5439 0.3673 0.4691 0.1273 0.3092 0.0502 0.0437 0.0750 0.0251 0.0264 0.1223 -
19 0.4602 0.4767 0.4790 0.1262 0.0892 0.0629 0.2598 0.0854 0.0199 0.0152 0.1196 -
Table 17: Maximum Persistence Score Using Excess of Mass (EoM) Criteria on Layer B3-POST
Table 18: Maximum Persistence Score Using Leaves of the Tree (Leaf) Criteria on Layer B3-POST
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 0.4276 0.5156 0.3370 0.0762 0.7069 0.6149 0.6534 0.0883 0.3577 0.2428 0.2445 -
5 0.2870 0.4289 0.4474 0.0576 0.3106 0.3816 0.4411 0.0456 0.2375 0.1646 0.3076 -
7 0.4534 0.5185 0.5102 0.0303 0.3663 0.2668 0.3402 0.0393 0.2188 0.2337 0.2005 -
9 0.3626 0.5545 0.5050 0.0227 0.3399 0.2858 0.3509 0.0365 0.3094 0.1886 0.3123 -
11 0.5077 0.5282 0.5285 0.0308 0.3258 0.3563 0.3273 0.0367 0.3043 0.3254 0.1462 -
13 0.4056 0.4398 0.5302 0.0348 0.3474 0.4060 0.2538 0.0344 0.2770 0.1902 0.1532 -
15 0.4491 0.4529 0.3906 0.0324 0.3106 0.3339 0.2409 0.0383 0.2639 0.2010 0.2788 -
17 0.3476 0.4070 0.4813 0.0318 0.2622 0.2003 0.2354 0.0367 0.2168 0.1933 0.1561 -
19 0.4228 0.4132 0.4711 0.0330 0.2438 0.2732 0.2147 0.0326 0.3260 0.2037 0.2560 -
Table 18: Maximum Persistence Score Using Leaves of the Tree (Leaf) Criteria on Layer B3-POST
Table 19: Mean Persistence Score Using Excess of Mass (EoM) Criteria on Layer B3-POST
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 0.0365 0.0416 0.0400 0.0232 0.0315 0.0196 0.0447 0.0243 0.0135 0.0076 0.0096 -
5 0.0333 0.0533 0.0318 0.0200 0.0508 0.0601 0.0562 0.0495 0.0211 0.0083 0.0065 -
7 0.0361 0.0406 0.0450 0.0150 0.0482 0.0582 0.0664 0.0390 0.0069 0.0138 0.0029 -
9 0.0564 0.0456 0.0381 0.0324 0.0580 0.0510 0.0646 0.0144 0.0136 0.0115 0.0088 -
11 0.0426 0.0527 0.0448 0.0310 0.0562 0.0563 0.0588 0.0208 0.0136 0.0314 0.0098 N/A
13 0.0534 0.0651 0.0538 0.0250 0.0592 0.0520 0.0817 0.0195 0.0270 0.0208 0.0091 -
15 0.0601 0.0484 0.0577 0.0319 0.0576 0.0526 0.0756 0.0240 0.0285 0.0111 0.0196 -
17 0.0592 0.0569 0.0556 0.0464 0.0530 0.0492 0.0385 0.0374 0.0193 0.0192 0.0209 -
19 0.0688 0.0665 0.0710 0.0347 0.0587 0.0526 0.0442 0.0402 0.0188 0.0150 0.0189 -
Table 19: Mean Persistence Score Using Excess of Mass (EoM) Criteria on Layer B3-POST
Table 20: Mean Persistence Score Using Leaves of the Tree (Leaf) Criteria on Layer B3-POST
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 0.0306 0.0277 0.0329 0.0091 0.0378 0.0251 0.0382 0.0158 0.0140 0.0128 0.0125 -
5 0.0320 0.0349 0.0295 0.0090 0.0417 0.0564 0.0559 0.0135 0.0137 0.0124 0.0097 -
7 0.0407 0.0384 0.0337 0.0078 0.0568 0.0508 0.0466 0.0143 0.0108 0.0122 0.0146 -
9 0.0349 0.0411 0.0482 0.0073 0.0567 0.0540 0.0542 0.0133 0.0132 0.0140 0.0109 -
11 0.0371 0.0431 0.0509 0.0079 0.0577 0.0556 0.0562 0.0138 0.0138 0.0199 0.0163 -
13 0.0482 0.0495 0.0534 0.0094 0.0554 0.0538 0.0558 0.0136 0.0171 0.0225 0.0178 -
15 0.0532 0.0511 0.0527 0.0089 0.0514 0.0563 0.0492 0.0138 0.0205 0.0198 0.0256 -
17 0.0560 0.0527 0.0595 0.0116 0.0524 0.0479 0.0520 0.0146 0.0202 0.0227 0.0184 -
19 0.0615 0.0471 0.0558 0.0111 0.0483 0.0513 0.0532 0.0142 0.0283 0.0230 0.0204 -
Table 20: Mean Persistence Score Using Leaves of the Tree (Leaf) Criteria on Layer B3-POST
Table 21: Number of Non-Clustered Samples Using Excess of Mass (EoM) Criteria on Layer OUT-POST
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 14662 14607 14815 21039 13825 13508 13837 30776 26055 27664 27759 40000
5 16383 16622 16725 21420 15423 15846 16184 31424 30657 30204 30662 40000
7 16880 17850 17518 22162 17062 17231 6129 33308 2701 29328 31931 40000
9 16617 17526 18508 22614 17226 17659 3 33089 23601 13626 23970 40000
11 16638 17407 18036 31452 12 16675 11718 33415 15867 20965 26013 40000
13 17593 16389 15931 25312 183 16959 35 34758 27182 29422 14139 40000
15 16800 17619 17985 24558 279 15081 17303 34567 27865 25974 22233 40000
17 15452 18595 15727 22344 16880 17777 0 34221 7434 4143 22767 40000
19 16563 18416 17879 22468 0 126 15206 33849 15351 24359 2312 40000
Table 21: Number of Non-Clustered Samples Using Excess of Mass (EoM) Criteria on Layer OUT-POST
Table 22: Number of Non-Clustered Samples Using Leaves of the Tree (Leaf) Criteria on Layer OUT-POST
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 18063 18541 18614 35151 17361 17395 17053 31801 28788 30572 30775 -
5 21315 21405 21396 37230 19935 19872 20022 35515 33103 33501 33810 -
7 21891 22380 22272 37596 20956 21143 21137 36435 33945 33905 33845 -
9 22346 22380 22862 38019 21555 21678 21661 36834 33954 33654 33760 -
11 23126 23334 23169 38306 21975 21998 22759 37121 34088 34178 33594 -
13 22558 23596 23210 38322 22326 22186 22310 36963 34472 33850 34476 -
15 22615 22867 22865 38399 22567 22177 22761 36492 33828 33555 34112 -
17 22701 23375 23341 37897 22639 22500 22584 36429 34182 33987 33922 -
19 22761 23588 23981 37875 22489 22119 22924 36357 34082 34075 34346 -
Table 22: Number of Non-Clustered Samples Using Leaves of the Tree (Leaf) Criteria on Layer OUT-POST
Table 23: Number of Clusters Using Excess of Mass (EoM) Criteria on Layer OUT-POST
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 2640 2642 2533 485 2711 2628 2737 1287 1694 1356 1345 0
5 1176 1191 1135 139 1293 1321 1331 296 473 358 375 0
7 758 768 755 85 833 863 410 146 2 169 180 0
9 543 531 573 45 613 625 2 107 73 36 72 0
11 419 430 446 48 2 455 285 83 33 38 71 0
13 358 323 313 28 3 373 2 74 70 70 32 0
15 287 306 308 21 2 287 314 52 61 55 54 0
17 237 279 212 16 251 297 2 49 3 2 49 0
19 219 242 233 15 2 2 211 42 26 45 2 0
Table 23: Number of Clusters Using Excess of Mass (EoM) Criteria on Layer OUT-POST
Table 24: Number of Clusters Using Leaves of the Tree (Leaf) Criteria on Layer OUT-POST
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 3176 3190 3117 876 3233 3249 3308 1457 1946 1591 1584 -
5 1526 1511 1498 259 1586 1625 1631 413 573 468 445 -
7 1001 1008 982 145 1067 1073 1067 204 283 258 234 -
9 727 711 718 89 785 789 774 132 177 164 161 -
11 572 572 553 67 635 615 621 96 131 129 129 -
13 470 489 465 54 528 526 525 78 104 101 103 -
15 416 398 400 37 444 444 442 59 94 97 97 -
17 352 362 364 30 397 384 402 52 84 81 86 -
19 314 319 321 28 342 340 353 44 78 82 76 -
Table 24: Number of Clusters Using Leaves of the Tree (Leaf) Criteria on Layer OUT-POST
Table 25: Maximum Persistence Score Using Excess of Mass (EoM) Criteria on Layer OUT-POST
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 0.3694 0.4264 0.4474 0.3099 0.3946 0.4431 0.6246 0.2298 0.4495 0.4024 0.3603 N/A
5 0.5160 0.3609 0.5841 0.0916 0.2771 0.2897 0.2785 0.0144 0.3540 0.3793 0.3057 N/A
7 0.5897 0.5506 0.5459 0.2107 0.5769 0.4094 0.3126 0.0239 0.0202 0.2876 0.3463 N/A
9 0.5327 0.5661 0.6323 0.2115 0.6997 0.4215 0.1101 0.0472 0.3223 0.2798 0.3862 N/A
11 0.4807 0.5102 0.5197 0.2359 0.0483 0.5354 0.5419 0.0511 0.2833 0.2636 0.3071 N/A
13 0.5894 0.5763 0.4898 0.1984 0.0676 0.4918 0.0844 0.0988 0.3292 0.2972 0.2758 N/A
15 0.5321 0.5367 0.4605 0.1571 0.0716 0.5754 0.5508 0.1044 0.3440 0.3016 0.2799 N/A
17 0.5079 0.4716 0.5201 0.1687 0.5797 0.5820 0.0920 0.1334 0.2208 0.2897 0.2789 N/A
19 0.4999 0.4819 0.5183 0.1832 0.1315 0.0510 0.5731 0.1437 0.2788 0.2857 0.0304 N/A
Table 25: Maximum Persistence Score Using Excess of Mass (EoM) Criteria on Layer OUT-POST
Table 26: Maximum Persistence Score Using Leaves of the Tree (Leaf) Criteria on Layer OUT-POST
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 0.4971 0.4754 0.5048 0.2039 0.3971 0.4208 0.5315 0.2347 0.4626 0.3048 0.3650 -
5 0.4111 0.6042 0.4260 0.3325 0.2472 0.3480 0.3081 0.1836 0.2701 0.2882 0.4539 -
7 0.5469 0.5549 0.4893 0.2627 0.6517 0.4215 0.4710 0.1405 0.2804 0.2992 0.2883 -
9 0.5988 0.5699 0.4852 0.2415 0.5191 0.3972 0.4925 0.1455 0.2607 0.2722 0.3869 -
11 0.4737 0.4981 0.5061 0.2149 0.4834 0.6152 0.6066 0.1525 0.3024 0.2890 0.2878 -
13 0.4875 0.6657 0.6903 0.1919 0.6927 0.3596 0.1544 0.1553 0.2727 0.3135 0.3068 -
15 0.5322 0.5847 0.5398 0.1798 0.5753 0.6147 0.7089 0.1667 0.3018 0.2986 0.2961 -
17 0.4606 0.4636 0.5500 0.1687 0.5912 0.5849 0.5581 0.1691 0.2911 0.3127 0.2777 -
19 0.4828 0.4422 0.4815 0.1832 0.6536 0.6201 0.6510 0.1740 0.2720 0.2809 0.3200 -
Table 26: Maximum Persistence Score Using Leaves of the Tree (Leaf) Criteria on Layer OUT-POST
Table 27: Mean Persistence Score Using Excess of Mass (EoM) Criteria on Layer OUT-POST
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 0.0281 0.0202 0.0082 0.0022 0.0141 0.0291 0.0406 0.0025 0.0218 0.0207 0.0154 N/A
5 0.0329 0.0237 0.0334 0.0015 0.0277 0.0270 0.0310 0.0005 0.0214 0.0306 0.0302 N/A
7 0.0374 0.0304 0.0374 0.0033 0.0407 0.0424 0.0376 0.0008 0.0105 0.0525 0.0472 N/A
9 0.0467 0.0442 0.0520 0.0053 0.0370 0.0395 0.0720 0.0009 0.0749 0.1184 0.1000 N/A
11 0.0503 0.0447 0.0571 0.0053 0.0358 0.0301 0.0358 0.0011 0.1090 0.1373 0.1073 N/A
13 0.0512 0.0551 0.0441 0.0076 0.0401 0.0318 0.0734 0.0019 0.0879 0.0967 0.1199 N/A
15 0.0623 0.0777 0.0535 0.0081 0.0466 0.0621 0.0304 0.0025 0.0881 0.1145 0.1263 N/A
17 0.0774 0.0759 0.0706 0.0112 0.0326 0.0540 0.0748 0.0031 0.1114 0.1579 0.1312 N/A
19 0.0788 0.0674 0.0635 0.0129 0.1231 0.0485 0.0545 0.0040 0.1049 0.1108 0.0247 N/A
Table 27: Mean Persistence Score Using Excess of Mass (EoM) Criteria on Layer OUT-POST
Table 28: Mean Persistence Score Using Leaves of the Tree (Leaf) Criteria on Layer OUT-POST
Aggregation Method Proposed Baseline No aggregation
UMAP Dimensionality 5 10 15 N/A 5 10 15 N/A 5 10 15 N/A
Min Cluster Size
3 0.0187 0.0165 0.0121 0.0015 0.0186 0.0229 0.0414 0.0023 0.0148 0.0149 0.0196 -
5 0.0190 0.0145 0.0224 0.0022 0.0246 0.0463 0.0333 0.0023 0.0266 0.0239 0.0277 -
7 0.0285 0.0334 0.0135 0.0034 0.0265 0.0465 0.0338 0.0024 0.0295 0.0384 0.0472 -
9 0.0380 0.0306 0.0360 0.0043 0.0302 0.0451 0.0409 0.0019 0.0368 0.0506 0.0561 -
11 0.0393 0.0480 0.0399 0.0046 0.0171 0.0347 0.0318 0.0028 0.0546 0.0648 0.0640 -
13 0.0435 0.0441 0.0394 0.0045 0.0395 0.0297 0.0270 0.0028 0.0591 0.0794 0.0693 -
15 0.0626 0.0631 0.0617 0.0056 0.0383 0.0397 0.0337 0.0033 0.0651 0.0696 0.0709 -
17 0.0623 0.0653 0.0701 0.0061 0.0210 0.0427 0.0339 0.0036 0.0619 0.0856 0.0779 -
19 0.0629 0.0574 0.0589 0.0070 0.0308 0.0365 0.0470 0.0045 0.0668 0.0753 0.0843 -
Table 28: Mean Persistence Score Using Leaves of the Tree (Leaf) Criteria on Layer OUT-POST